OpenStudy (anonymous):

Why is 1 m^3 = 1 000 000 ml?

OpenStudy (anonymous):

would it help if i gave you the conversion factor 1cm^3=1 ml

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

like how did 1.0 m^3 get to 1 000 000 ml..o.o

OpenStudy (anonymous):

\[1 M^3(\frac{100 CM^3}{1 M^3})(\frac{1 ML}{1 CM^3})=1,000,000 ML\] Remember you have to multiply 100 x 100 x 100 which equals 1 million

OpenStudy (anonymous):

so when I cancel m3, and cm 3... 1 x 100 x 1 = 1 000 000 mL? o.O

OpenStudy (anonymous):

when you have something raised to a power in a factor lable thing you have to repeat that step as many times as the power says. in this case you do it 3 times. So for the first part you get 1,000,000 Cm^3 then cancel out the Cm^3 and leave you with 1,000,000 ML

OpenStudy (anonymous):

ohhh.. but isn't that cubic cm?

OpenStudy (anonymous):

Yes but the cubic cm cancels out because you have one on the top of the fraction and one on the bottom of the fraction. And cubic meters cancel out because you have one on top and one on bottom and this only leaves you with ML

OpenStudy (anonymous):

ohh.. i see, it's just 'cause I saw like a similar question where the askeer asked: "Why 1 m^3 = 1000 L" and a person answered with: 1 m^3 * (100 cm / 1 m)^3 * (1 mL / 1 cm^3) * (1 L / 1000 mL) = 100^3 / 1000 = (10^2)^3 / 10^3 = 10^6 / 10^3 = 1000 L

OpenStudy (anonymous):

You got it

OpenStudy (anonymous):

ahh, i still don't get it D: